OpenStudy (rubioandrea1649):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations?

jhonyy9 (jhonyy9):

for example : 2x +3y =6 multiplie by 2 5x +2y =4 multiplie by -3 4x +6y =12 -15x -6y = -12 ---------------- -11x 0 = 0 hope this will help you